# Interpreting the Hi cm cosmology maps through Largest Cluster Statistics. Part II. Impact of the realistic foreground and instrumental noise on synthetic SKA1-Low observations

### v1

@@ 5.1 Effect of thresholding
This work presents an optimum strategy to reduce the bias in LCS estimation from the simulated 21 cm observation results. Our previous study, Dasgupta et al. [33] used a thresholding method, e.g., gradient descent, to binarize 21 cm maps and estimate the LCS. During the early stages of reionization, the histogram of 21 cm maps is not bimodal. Instead, it exhibits an unimodal distribution with asymmetric tails, as indicated by non-zero skewness [34, 66]. This thresholding method introduces bias for these cases due to its limitations in reaching the local minima in the 21 cm observation maps. The problem becomes more significant when telescope effects, such as the proper implementation of the array synthesized beam of SKA1-Low, are included. Therefore, these limitations fail to binarize the image pixels and lead to biases in the recovered percolation process of reionization. In order to mitigate these challenges, we use the triangle method to separate the neutral and ionized regions. This method identifies the threshold by drawing a line from the peak of the histogram to the end of the asymmetric tail and determining the point of maximum distance between this line and the histogram. The visual representation of the triangle method is shown in Figure 5. This method effectively separates ionized and neutral regions during different stages of EoR from the noiseless data [67]. In Figure 6, we compare the estimation of LCS obtained from the simulated 21 cm observation against neutral fraction using different thresholding methods. The original image is a hypothetical scenario without telescope effect and noise, and the corresponding threshold for LCS is set at zero. The red dash-dot and green dashed curves represent the estimation of LCS based on the thresholds obtained via the triangle and gradient-descent methods, respectively. The triangle method demonstrates superior performance compared to the gradient-descent method. We observe that the threshold set by the triangle method on the simulated 21 cm observations results in the same percolation transition redshift or [34, 36, 29, 32, 33] as that of the hypothetical scenario, i.e. without any telescope effect and noise shown by the blue solid curve. However, the obtained LCS based on the threshold set by the triangle method has small deviations from the original LCS estimation, which is due to the resolution of the telescope and the error in the thresholding method.
@@ 5.2 Effect of gain calibration error
This section investigates the impact of antenna-based gain calibration errors on the obtained LCS from synthetic 21 cm observations with SKA1-Low. The analysis is based on the threshold set via the triangle method on residual image cubes. There are many techniques that have been developed to remove residual contamination from the image domain [4]. However, we aim to determine the maximum tolerance level of antenna-based calibration error to recover the reionization history in a robust and relatively unbiased manner using LCS. We varied antenna-based calibration errors to estimate this tolerance level to achieve a sufficient dynamic range to extract the 21 cm signal from residual data. Here, two types of foreground corruption conditions were considered, as discussed in subsection 2.2. We discuss the results obtained for the different cases of foreground models below.
@@
Case I: Figure 7 shows the evolution of obtained LCS against for varying levels of calibration errors in the sky models for Case I. These residual data consist of residual point sources and the target signal. The blue solid curve represents the original image, a hypothetical scenario without telescope effect and noise, and the corresponding threshold for LCS is set at zero. The magenta circle, cyan diamond, and green square curves represent the obtained LCS from simulated 21 cm observation via 21cmE2E pipeline based on the threshold set by the triangle method for different post-gain calibration errors of , , and , respectively. We observed that obtained LCS from simulated 21 cm observations with calibration errors up to follow with the obtained LCS without any bias due to corruption, as shown by the red star curve. However, beyond a calibration inaccuracy of , the computation of LCS becomes challenging. At this level, RMS noise overwhelms the image pixels, and the imperfect removal of foreground residuals significantly impacts the bimodal distribution of simulated 21 cm maps. This corruption makes it challenging to separate ionized and neutral regions accurately. At higher calibration errors, this thresholding method fails to recover ionized regions due to biases in the histogram of the bimodal distribution. This thresholding method misidentified unsubtracted residual contamination as neutral regions and other regions as ionized regions, as shown in the bottom-right panel of Figure 3. Therefore, the SURFGEN2 code provides the value of LCS close to unity, as indicated by the green square line in Figure 7.
Case II: Figure 8 shows the evolution of LCS against for different levels of calibration errors in the sky models for Case II. The sky models consist of residual point sources, diffuse emission, and the target signal. The magenta circle, cyan diamond, and green square curves represent the obtained LCS from simulated 21 cm observation via 21cmE2E pipeline based on the threshold set by the triangle method for different post-gain calibration errors of , , and , respectively. We observed that obtained LCS from simulated 21 cm observations with calibration errors up to follow with the obtained LCS without any bias due to corruption, as shown by the red star curve. However, with calibration inaccuracy of , the computation of LCS becomes challenging as residual foreground contamination starts to dominate the image and introduces bias in LCS analysis. A similar effect of gain calibration errors on the obtained LCS from simulated 21 cm observations is observed across both groups of foreground models.
@@
We observed that calibration errors introduce artificial filamentary or tunnel-like features (deconvolution artefacts) in the final image cube. These artefacts fragment the largest ionized region into isolated regions. The fragmentation depends on the thresholding method used to separate the Hi field. The left panel of Figure 8 highlights the impact of these deconvolution artefacts, such as artificial filamentary or tunnel-like features, on the LCS analysis, particularly in the dip scenario. It is essential to note that imperfect subtraction of bright sources can also create a negative bowl-like region. These regions further fragment the largest ionized region into pieces, introducing additional bias in the LCS analysis. Our initial systematic study emphasizes the critical importance of achieving calibration accuracy for the upcoming SKA1-Low observations. Such accuracy is essential to recover the reionization history in a robust and unbiased manner using the LCS. However, the choice of thresholding algorithm significantly impacts the LCS analysis. Future work will explore different thresholding algorithms to improve the separation of neutral and ionized regions during LCS estimation.
@@ 5.3 Effect of instrumental noise
Figure 10 shows the evolution of obtained LCS from 21 cm observation against for different observation times and imaging weighting. For deep 21 cm observations with 1500 hours and 2000 hours of integration time, the obtained LCS up to , threshold set by Yen’s method shown by magenta square and black cross curves, respectively consistent with the estimation of LCS from the original downsampled image cube, represented by a red dotted curve. The choice of weighting scheme significantly affects the properties of synthesized image cubes. With the natural weighting scheme, the resolution of the synthesized image cubes is relatively lower than that of the briggs weighting scheme. However, it offers a higher signal-to-noise ratio (SNR) in the final maps than briggs weighting. After adding instrumental noise, small-scale filamentary structures appear in the image cubes. In the natural weighting scheme, larger ionized regions contribute to LCS estimation due to lower resolution and higher SNR. In contrast, in the briggs weighting scheme, the poorer SNR causes numerous small filamentary structures that may be classified as either ionized or neutral, depending on the thresholding algorithm. Thus, the choice of the weighting scheme affects the synthesized maps, which in turn impacts the obtained LCS from the 21 cm observation. We observed that the features of the obtained LCS could be computed up to a certain neutral fraction , and the percolation transition point shifts significantly toward lower redshifts. This shift is caused by the combined effects of the instrumental noise and the partially ionized regions in the Hi21 cm field. Thus, this leads to a biased interpretation of the history of reionization. Our findings highlight the significance of imaging weighting to obtain the LCS from 21 cm observation. In addition, it emphasized the importance of the thresholding algorithm to separate neutral and ionized pixels in noisy data and the impact of cluster detection within the SURFGEN2 code on LCS estimation. Future work will explore alternative thresholding algorithms for improved pixel differentiation.
@@ 6 Summary and discussion
The histogram of the brightness temperature map of the 21 cm field during EoR exhibits a bimodal distribution. The two peaks correspond to ionized and neutral regions. In a hypothetical scenario without telescope effects or noise, ionized regions would have a brightness temperature of zero. However, in radio interferometric observation, determining an optimal threshold to distinguish neutral from ionized regions in the 21 cm brightness map is challenging. The complexity arises from the systematic effects, resolution of the telescope, and thermal noise, which cause random shifts in pixel brightness temperature maps. Therefore, accurately differentiating between these regions in 21 cm maps becomes difficult. In this study, we present alternative thresholding strategies to binarize the image pixels to recover the percolation transition. It was observed that the bias of estimation of LCS of the 21 cm maps reduces after employing the triangle method of thresholding. Furthermore, we observed that the threshold set by the triangle method on the simulated 21 cm observations results in the same percolation transition redshift or as a hypothetical scenario, i.e. without any telescope effect and noise. However, the obtained LCS has a little deviation from the actual LCS estimation from the hypothetical scenario due to the resolution of the SKA1-Low telescope.
@@
We study the effect of direction-independent calibration errors on the percolation process of reionization history using the LCS of the 21 cm maps. We have demonstrated that a post-calibration and post-averaging antenna-based calibration error tolerance of is essential to achieve unbiased and unaffected LCS estimation. This corresponds to an amplitude error of and a phase error of in each time domain. It is important to note that the tolerance of will vary depending on the RMS variation across different sky patches, i.e., the statistical distribution of bright sources within the FoV. We also observed that the threshold set by the triangle method on the residual 21 cm maps simulated via 21cmE2E aligns with the same percolation transition redshift or , as results from the simulated 21 cm observations without any bias due to corruption. However, at higher calibration errors( ), the dynamic range of the 21 cm maps deteriorates. The RMS noise from the residual foreground starts to dominate the image pixels and affects the sharp features of the bimodal distribution of the Hi cm field. Under these conditions, the thresholding method, e.g. the triangle method, struggles to distinguish neutral and ionized regions. This results in a biased estimation of LCS.
The calibration errors introduce artefacts, such as artificial filament structures, in the observed Hi maps during deconvolution. When the thresholding algorithm is used to binarize the image pixels, these artefacts cause the largest ionized region to fragment into isolated segments, thereby introducing dips in the LCS evolution for low . The peeling of bright sources can further break down the largest ionized regions. This fragmentation depends on the spatial distribution of the point sources.
Next, we introduce the instrumental noise for SKA1-Low to assess the robustness of the evolution of the obtained LCS. After introducing the instrumental noise, it is observed that the bimodal feature of the 21 cm map is lost, and the histogram shifts towards a Gaussian distribution. Our initial analysis used the triangle method to binarize the image pixels. However, it later proved to be inadequate for determining an optimal threshold, resulting in inaccurate percolation analysis during reionization. To address this, we used an entropy-based thresholding algorithm, Yen’s method. This method determined the optimal threshold for minimizing the entropy of the pixel intensity distribution. We observed that with an accumulation of h of observation time, the estimation of the LCS threshold set by Yen’s method remains unchanged and unbiased from the results of the estimation of LCS from the original downsampled image cubes. This accumulation of observation time represents an optimistic test for the robustness of LCS under realistic conditions. However, during the early stages of reionization, i.e., at the higher neutral fractions, the feature of LCS is lost owing to random fluctuations in the thermal noise that dominate the ionization fluctuations. Furthermore, for total observation periods smaller than hours, the lower SNR significantly limits the computation of the estimation LCS from 21 cm observation. As the contrast between ionized bubbles and random fluctuations becomes comparable, the performance of the LCS analysis is impacted.
@@
We demonstrate that the obtained LCS results can be biased depending on the choice of thresholding method. Therefore, the development of a robust thresholding method is essential for precise percolation analysis across all reionization stages. However, we should have accounted for the impact of ionospheric phase distortion, the chromatic nature of the primary beam, and other systematic problems. These factors could further complicate the LCS analysis. Our future work will address these factors to provide a more comprehensive understanding of their implications for EoR cm, and subsequent LCS analysis.

### v2

@@ 5.1 Effect of thresholding
This work presents an optimum strategy to reduce the bias in LCS estimation from the simulated 21 cm observation results. Our previous study, Dasgupta et al. [33] used a thresholding method, e.g., gradient descent, to binarize 21 cm maps and estimate the LCS. During the early stages of reionization, the histogram of 21 cm maps is not bimodal. Instead, it exhibits an unimodal distribution with asymmetric tails, as indicated by non-zero skewness [34, 67]. This thresholding method introduces bias for these cases due to its limitations in reaching the local minima in the 21 cm observation maps. The problem becomes more significant when telescope effects, such as the proper implementation of the array synthesized beam of SKA1-Low, are included. Therefore, these limitations fail to binarize the image pixels and lead to biases in the recovered percolation process of reionization.
In order to mitigate these issues, we propose an optimum thresholding method for robustly separating neutral and ionized regions. In figure 5, we plot the comparison of the obtained LCS from the simulated 21 cm observation against neutral fraction using different thresholding methods. The original image is a hypothetical scenario without telescope effect and noise, and the corresponding threshold for LCS is set at zero. The red dash-dot and green dashed curves represent the estimation of LCS based on the thresholds obtained via the optimum thresholding method and gradient-descent methods, respectively. The optimum thresholding method demonstrates superior performance compared to the gradient-descent method. We observe that the threshold set by the optimum thresholding method on the simulated 21 cm observations results in the same percolation transition redshift or [34, 36, 29, 32, 33] as that of the hypothetical scenario, i.e. without any telescope effect and noise shown by the blue solid curve. However, the obtained LCS based on the threshold set by the optimum thresholding method has deviations from the original LCS estimation, due to the resolution of the telescope and the error in the thresholding method.
@@ 5.2 Effect of gain calibration error
In this section, we discuss the impact of time-frequency correlated antenna-based residual gain calibration errors on the obtained LCS from 21 cm observation results. This LCS analysis is based on the threshold set via the optimum thresholding method on residual image cubes. There are many techniques that have been developed in the past decade to remove residual contamination from the image domain [4]. However, we aim to determine the maximum tolerance level of antenna-based calibration error to recover the reionization history in a robust and relatively unbiased manner using LCS. In order to estimate the tolerance level, we increase the level of calibration errors. This approach helps us ascertain the dynamic range required to extract the 21 cm signal effectively from the residual data. For this work, we consider two types of foreground contamination in our LCS analysis, as detailed in Section 2.2. The results obtained for these different foreground models are discussed in the subsequent subsections.
@@
Case I: Figure 6 compares the evolution of the obtained LCS with for varying levels of residual calibration errors on the simulated SKA1-Low observational maps. These maps include both unsubtracted point source contamination and the target 21-cm signal. The effect of residual calibration errors on LCS estimation is shown in magenta circle, cyan diamond, and green square curves in figure 6. These results are obtained from the simulated 21 cm observation via 21cmE2E pipeline based on the threshold set by the optimum thresholding method. The blue solid curve represents the original image, a hypothetical scenario without telescope effect and noise, and the corresponding threshold for LCS is set at zero. We observed that at the residual gain calibration, the inaccuracy of (illustrated by the red star curve), the obtained LCS from simulated 21 cm observations remains largely unaffected. It closely matches the evolution of the LCS as seen in the obtained LCS from 21 cm observations without any bias due to corruption.
However, with higher calibration errors, the computation of LCS from the simulated SKA1-Low observational maps becomes challenging. This occurs because the residual foreground, resulting from residual gain calibration errors, introduces artificial filamentary or tunnel-like features (deconvolution artefacts) in the final image cube. These artefacts lead to the fragmentation of the largest ionized region into isolated regions. The extent of this fragmentation depends on the thresholding method used to binarize each Hi map. Furthermore, this systematic effect reduces the contrast between ionized and neutral pixels. This reduction is primarily due to the additional RMS noise introduced by the residual contamination, which in turn lowers the overall signal-to-noise ratio (SNR). Therefore, this contamination makes it challenging to identify the ionized and neutral regions accurately and leads to a biased interpretation of the history of reionization using LCS.
Case II: Figure 7 compares the evolution of the obtained LCS with for different levels of residual calibration errors on the simulated SKA1-Low observational maps for Case II. These maps include unsubtracted point source contamination, diffuse emission and the target signal. The magenta circle, cyan diamond, and green square curves illustrate the effect of the residual post-calibration errors on LCS estimation in figure 7. Similar to Case I, we also observed with the residual gain calibration, the inaccuracy of (illustrated by the red star curve), the obtained LCS from simulated 21 cm observations remains largely unaffected. It closely matches the evolution of the LCS as seen in the obtained LCS from 21 cm observations without any bias due to corruption. Therefore, a post-calibration inaccuracy of is required to recover the history of reionization in a robust and relatively unbiased manner using LCS. Although a calibration error of , the apparent percolation curves visually appear to match with a zero-error case. However, this small error can still substantially contribute to image artefacts and morphological distortions.
@@
It is essential to note that imperfect subtraction of bright sources can also create negative bowl-like regions. These regions further fragment the largest ionized regions into isolated pieces, thereby introducing additional bias into the LCS analysis. Our initial systematic study emphasizes the critical importance of achieving high calibration accuracy for the upcoming SKA1-Low observations. Such accuracy is essential to recover the reionization history in a robust and unbiased manner using the LCS. However, the choice of thresholding algorithm significantly impacts the LCS analysis. Future work will explore various thresholding algorithms to improve the separation of neutral and ionized regions during LCS estimation.
@@ 5.3 Effect of instrumental noise
In this section, we discuss the effect of the instrumental noise using a realistic simulation on the synthesized SKA1-Low observational maps. In section 3.3, we discuss the prescription to add instrument noise in the visibility domain. Figure 9 illustrates the evolution of the obtained LCS from 21 cm observation against for different deep observation times. The original image is a hypothetical scenario without telescope effect and noise, and the corresponding threshold for LCS is set at zero, indicated by a blue solid curve. A green dotted curve illustrates the obtained LCS from the observed downsample image, based on the threshold set by the optimum thresholding method. For deep 21 cm observations with 1500 hours and hours of integration time, the obtained LCS (threshold set by the optimum thresholding method) is shown by the magenta and black curves, respectively. It is observed that, although the LCS features can be computed up to a certain , the percolation transition threshold systematically shifts towards a lower . This systematic shift arises from the combined effects of random noise fluctuations and the introduction of partially ionized regions within the HI 21 cm field. Consequently, such a systematic bias in the percolation transition threshold may lead to a significantly biased interpretation of the reionization history. This similar feature of LCS is observed when considering the combined effects of additive Gaussian noise in the image plane and the choice of larger smoothing scales [33].
Our findings highlight the significance of imaging weighting to obtain the LCS from 21 cm observation. In addition, it emphasized the importance of the thresholding algorithm to separate neutral and ionized pixels in noisy data and the impact of cluster detection within the SURFGEN2 code on LCS estimation. Future work will explore alternative thresholding algorithms for improved pixel differentiation.
@@ 6 Summary and discussion
The histogram of the brightness temperature map of the 21 cm field during EoR exhibits a bimodal distribution. The two peaks correspond to ionized and neutral regions. In a hypothetical scenario without telescope effects or noise, ionized regions would have a brightness temperature of zero. However, in radio interferometric observation, determining an optimal threshold to distinguish neutral from ionized regions in the 21 cm brightness map is challenging. The complexity arises from the systematic effects, resolution of the telescope, and thermal noise, which cause random shifts in pixel brightness temperature maps. Therefore, accurately differentiating between these regions in 21 cm maps becomes difficult. In this study, we introduce an optimal thresholding strategy to binarize the image pixels and accurately recover the percolation transition. It is observed that this method significantly reduces the bias in LCS estimation from 21cm maps. However, the obtained LCS has little deviation from the actual LCS estimation from the hypothetical scenario due to the resolution of the SKA1-Low telescope and error in the thresholding method.
@@
We study the effect of direction-independent calibration errors on the percolation process of reionization history using the LCS of the 21 cm maps. We have demonstrated that a post-calibration and post-averaging antenna-based calibration error tolerance of is essential to achieve unbiased and unaffected LCS estimation. This corresponds to an amplitude error of and a phase error of in each time domain. It is important to note that the tolerance of will vary depending on the RMS variation across different sky patches, i.e., the statistical distribution of bright sources within the FoV. We also observed that the threshold set by the optimum thresholding method on the residual 21 cm maps simulated via 21cmE2E aligns with the same percolation transition redshift or , as results from the simulated 21 cm observations without any bias due to corruption. However, at higher calibration errors( ), the dynamic range of the 21 cm maps significantly deteriorates. The RMS noise from the residual foreground starts to dominate the image pixels and introduces spurious features that can either mimic or entirely obscure the faint 21 cm cosmological signal. Under these conditions, the thresholding method, fails to identify the boundary between the ionized and neutral pixels. This results in a biased estimation of LCS at the higher calibration errors.
Next, we introduce the instrumental noise for SKA1-Low to assess the robustness of the evolution of the obtained LCS. After introducing the instrumental noise, it is observed that the bimodal feature of the 21 cm map is lost, and the histogram shifts towards a Gaussian distribution. We observed that with an accumulation of h of observation time, the estimation of the LCS threshold set by the optimum thresholding method, systematically shifts towards a lower . This systematic shift arises from the combined effects of random noise fluctuations and the introduction of partially ionized regions within the HI 21 cm field. In the early stages of the EoR (i.e., at higher neutral fractions), the limited resolution causes small ionized regions to be confused with partially neutral within the Hi 21 cm field. Therefore, at this stage, the estimation of LCS could not be computed. As the contrast between ionized bubbles and random fluctuations becomes comparable, the performance of the LCS analysis is affected. At the later stage of reionization, the combination of poor resolution and instrumental noise leads to a biased determination of the threshold between ionized and neutral regions. This, in turn, results in erratic behaviour in the evolution of the recovered LCS from 21 cm observations and biased interpretation of reionization history. This feature of LCS is observed when considering the combined effects of additive Gaussian noise in the image plane and the choice of larger smoothing scales [33].
@@
Our findings demonstrate that the obtained LCS results can be significantly biased by the choice of thresholding method. This highlights the critical need for developing a robust thresholding technique to ensure precise percolation analysis across all reionization stages. However, we acknowledge that our current study did not account for the impact of ionospheric phase distortion, the chromaticity of the primary beam, and other systematic instrumental effects. These factors further complicate LCS analysis and potentially introduce additional biases. Our future work will address these factors to provide a more comprehensive understanding of their implications for EoR cm observations, and subsequent LCS analysis.
@@ 7 Acknowledgements
Appendix A Image Binarization
Threshold selection for identifying neutral and ionized pixels is a crucial aspect of EoR 21 cm image analysis. This work introduces a modified method that combines a global thresholding strategy with an unsharp masking technique, to identify ionized regions within 21cm observation maps. As illustrated in Figure [Figure Number], we compare our method with the binarization technique described by Giri et al. [66]. We find that, whereas their method tends to over-segment ionized regions, particully in areas with low-density fluctuations. Our method demonstrates superior robustness across diverse density environments and more accurately preserves the morphological structure of ionized regions. The black (white) regions in binary maps represent the ionized and neutral regions.